OpenStudy (bhuvan):

If Tan A -Tan B =x and Cot B - Cot A=y; Prove that:Cot (A-B)=1/x + 1/y

OpenStudy (mww):

Try this: cot(A-B) = 1/tan(A-B) and continue fro there.

OpenStudy (bhuvan):

Give full solution

OpenStudy (mjdennis):

From the Open Study Code of Conduct ( http://openstudy.com/code-of-conduct): OpenStudy values the Learning process - not the ‘Give you an answer’ process. @Bhuvan, most people will help you if you show that you are helping with your own learning. Most people won't just hand you answers.

OpenStudy (mww):

Recall tan(A-B) can be expanded \[LHS = \cot(A-B) = \frac{ 1 }{ \tan(A-B) } = \frac{ 1 }{ \frac{ \tan A - \tan B }{ 1+\tan A \tan B } }=\frac{ 1+\tan A \tan B }{ \tan A-\tan B }\] Then you sub in x and y into RHS 1/x and 1/y and compare.

OpenStudy (agent0smith):

"Give full solution" lol